# Generative Anchored Fields: Controlled Data Generation via Emergent Velocity Fields and Transport Algebra

### v1

@@ 5.3 Single-Class Generation (One $J$ and One $K$)
We first validate GAF on single-class generation, using the school bus class from ImageNet, which contains about 1.3k images used for training. Figure 8 shows some example generations. We can see that a single K head model with 250-step Heun sampling demonstrates high-quality, diverse generations. Note that the generations showcase varied viewpoints, lighting conditions, and backgrounds while maintaining class consistency.
@@ 5.4 Multi-Class Generation (One $J$ and Multiple $K$s)
We evaluate GAF on multi-class generation, using ImageNet-1k with $10$-class subsets and AFHQ ($512\times 512$ px) with three classes: cat, dog, and wild animals. Figure 9 and Figure 10 shows samples from each independent $K$ head, with each head corresponding to one of the ten or three classes with in each dataset. This demonstrates that GAF scales to multi-class generation while maintaining per-class quality. Each $K$ head produces diverse, high-fidelity samples within its class without interference from other heads.
Figure 9: Comprehensive gallery of generated samples from the $10$-class ImageNet subset. The figure demonstrates the model’s ability to handle high intra-class variance.
@@ 5.5.1 Pairwise Interpolation
We demonstrate direct $K$ head interpolation, by performing the operation $v=(1-\alpha)v_{i}+\alpha v_{j}$, for $\alpha\in[0,1].$ Figure 11 shows samples from interpolating between two classes from the AFHQ dataset, trained for multi-class generation in Section 5.4. We can see a smooth semantic transition between all three class pairs. Each interpolation produces coherent intermediate states, validating that independent $K$ heads enable linear composition in velocity space.
@@
Figure 11: Pairwise interpolation samples ($512\times 512$ px) from GAF trained with one $J$ and three independent $K$ heads on the AFHQ dataset (i.e., the same model as sampled in Figure 10). Each row shows an interpolation along a different transport direction. Note that we did not explicitly train GAF for interpolation or translation. Instead, GAF enables such operations inherently using Transport Algebra.
5.5.2 Continuous Cyclic Transport
[…]
We observe that all 3000 samples achieve LPIPS=0.0. This indicates GAF’s transport algebra is deterministically reversible. The perfect closure validates that independent $K$ heads form a consistent algebraic structure where cyclic compositions return exactly to their starting point.
@@ 5.5.2 Continuous Cyclic Transport
[…]
Figure 12: Continuous cyclic transport preserving identity across three classes. Each row shows a complete cycle starting from one class (left) to two interpolated classes, and finally returning to the initial class. Distinctive features (coloring, facial marking) are semantically preserved throughout each cycle while class-specific structure transforms.
@@ 5.5.3 Barycentric Transport
We extend pairwise interpolation to a three-way composition using barycentric coordinates. The blended velocity field is $v_{\text{blend}}=\alpha v_{i}+\beta v_{j}+\gamma v_{k}$, where $\alpha+\beta+\gamma=1$. We use square-to-simplex mapping with $\alpha=u(1-u),~\beta=(1-u)(1-v),$ and $\gamma=v$ to cover the full compositional space. Figure 13 shows three examples of three-way compositions, namely dog-cat-wild, cat-dog-wild, and wild-dog-cat. It shows all weighted combinations, with corners representing pure classes and the interior points showing multi-class blends.
@@ 6 Conclusion
We introduced Generative Anchored Fields (GAF), a model that learns independent endpoint predictors $J$ and $K$ rather than a direct trajectory. The velocity field v=KJv=K-J emerges from the time-conditioned disagreement between these endpoints. This factorization enables Transport Algebra, allowing compositional control through vector arithmetic on learned $K$ heads. We demonstrated that GAF achieves strong sample quality while uniquely providing lossless cyclic transport (LPIPS=0.0) and precise semantic manipulation as intrinsic architectural primitives.

### v2

@@ 5.2 Single-Class Generation (One $J$ and One $K$)
This section validates GAF on single-class generation, using the school bus class from ImageNet, which contains about 1.3k images used for training. Figure 7 shows some example generations. We can see that a single K head model with 250-step Heun sampling demonstrates high-quality, diverse generations. Note that the generations showcase varied viewpoints, lighting conditions, and backgrounds while maintaining class consistency.
@@ 5.3 Multi-Class Generation (One $J$ and Multiple $K$s)
We evaluate GAF on multi-class generation on three datasets. First, we utilize CIFAR-$10$ ($32\times 32$ px) with ten object classes and one $J$ head and ten $K$ heads. Second, we utilize AFHQ ($512\times 512$ px) with three classes: cat, dog, and wild animals, using one $J$ head and three $K$ heads. Third, we utilize ImageNet-1k with $1,000$ classes, using one $J$ head and $1,000$ $K$ heads. Figure 8 and Figure 9 show samples from each independent $K$ head, each corresponding to one of the three or thousand classes with in each dataset, respectively. This demonstrates that GAF scales to multi-class generation while maintaining high per-class quality. Each $K$ head produces diverse, high-fidelity samples within its class without interference from other heads.
Figure 9: Comprehensive gallery of generated samples from the $1,000$ class ImageNet dataset. The figure demonstrates the model’s ability to handle high intra-class variance.
5.4 Qualitative Evaluation
We analyze the effect of the number of IER, Euler, and Heun steps on generation quality using CelebA-HQ ($256\times 256$) and ImageNet ($256\times 256$). We quantitatively measure sample quality with the Fréchet Inception Distance (FID) [12, 37] and inception score (IS) [35], and the diversity using the recall score [19]. All FID scores are computed by the PyTorch Fidelity library [29] using 50,000 generated samples.
Figure 10: Effect of Euler steps on CelebA-HQ $256$ sample quality.
Figure 11: Effect of IER steps on CelebA-HQ $256$ sample quality.
As shown in Figure 10, with very few Euler steps (151-5) the samples are blurry but roughly coherent; increasing the steps to 5405-40 significantly improves the quality and semantic detail. The FID (calculated using Heun, Figure 12a) value decreases rapidly up to about $40$ steps, reflecting improved sample coherence, and then flattens, with strong diminishing returns beyond roughly $80$ steps. Beyond $40$ steps, the curve exhibits clear diminishing returns: FID improves from about $7.45$ at $40$ steps to only $7.27$ at $80$ steps. Therefore, while $80$ steps yield the best score, high-quality generation is already achieved around 2020-$40$ Heun steps. The IER sampler, of which samples are presented in Figure 11, converges with far fewer steps: requiring only 686-8 steps for high-quality generation; and additional IER steps provide diminishing results.
Table 3 and 4 reports FID scores on ImageNet $256$, CIFAR-10, and CelebA-HQ $256$ using the Heun sampler. On ImageNet $256$, GAF achieves an FID of $7.51$ (at $N=250$ steps), outperforming DiT-XL/2 ($9.6$) and SiT-XL ($8.3$), with competitive IS, precision and recall. On CelebA-HQ $256$, GAF achieves an FID of $7.27$ (at $N=80$ steps), trailing LDM-4 ($5.11$) which employs heavily optimized diffusion pipeline. GAF achieves an FID of $9.53$ on CIFAR-10 dataset.
Table 3: FID results on ImageNet $256\times 256$. All results without classifier-free guidance.
ImageNet 256
@@ 5.4 Qualitative Evaluation
DiT-XL/2 675 7M 9.6 121.50 0.67 0.67
SiT-XL/2 675 7M 8.3 131.65 0.68 0.67
GAF (ours, retrunked) 694 100k 7.51 114.81 0.54 0.69
Table 4: FID comparison on CelebeA-HQ $256\times 256$ and CIFAR-$10$
CelebA-HQ 256
[…]
Glow [17] 68.93
SDE [43] 7.23
LSGM [44] 7.22
LDM-4 [32] 5.11
GAF (ours) 7.27
CIFAR-10
[…]
DDPM 3.17
Score SDE 2.20
Flow Match. 6.35
GAF (ours) 9.53
Figure 12: Effect of the number of IER, and Heun steps on sample quality, measured by FID.
5.5 Ablation Study
The ablation study compares the full GAF objective, $\mathcal{L}_{\text{GAF}}=\mathcal{L}_{\text{pair}}+\lambda_{\text{res}}\,\mathcal{L}_{\text{res}}+\lambda_{\text{swap}}\,\mathcal{L}_{\text{swap}}$, against a pair-only variant trained with $\mathcal{L}_{\text{GAF}}=\mathcal{L}_{\text{pair}}$ alone (i.e, removing residual and swap losses).
5.5.1 Effect of Swap and Residual Loss
We report Pixel MSE, LPIPS(Alex), LPIPS(VGG), Latent MSE, and Latent LPIPS-GAF (computed from trunk features); lower is better for all metrics. The MSE captures the reconstruction error between the initial and final image, while LPIPS captures perceptual similarity. Latent space metrics measure transport consistency before VAE decoding.
Figure 13: Effect of swap and residual losses under cyclic transport ($z_{y}\rightarrow K_{0}\xrightarrow{J}K_{1}\xrightarrow{J}K_{2}\xrightarrow{J}K_{0}$).
As shown in Figure 13, across the cyclic transport, the full objective yields consistently low error rates and more stable multi-class transport at lower step counts. Removing $\mathcal{L}_{\text{res}}$ and $\lambda_{\text{swap}}$ does not collapse generation indicating that the $\mathcal{L}_{\text{pair}}$ loss is the dominant learning signal and the residual, and the time antisymmetric losses both serve as regulizers. At larger step counts ($N>100$), the error values of the full and pair objectives values converge. The $\mathcal{L}_{\text{res}}$ and $\lambda_{\text{swap}}$ are not required for endpoint learning; instead, they primarily act as stability and efficiency regularizers.
5.5.2 Image-to-Image Translation
Figure 14a shows an 8-step IER image-to-image translation using cross-modality transport between the cat and dog classes, as described in Section 3.3.2. The base images in Figure 14b are from the AFHQ held-out dataset. The images generated with GAF trained using the $\mathcal{L}_{\text{pair}}$-only objective (Figure 14c) show more drift and local artifact and inconsistent multi-class generation, while the full loss objective (Figure 14d) preserves identity and semantic structure more reliably under chained transport. This indicates that the residual and the swap losses are important for stabilization of the latent space.
Figure 14: Image-to-Image translation (8-step IER). (a) Visualization of the full cross-modality translation path ($K_{\text{cat}}\xrightarrow{J}K_{\text{dog}}$). (b) Base images from the held-out AFHQ dataset. (c) Translations generated by GAF trained only with pair loss ($\mathcal{L}_{\text{pair}}$), which exhibit significant drift and local artifacts. (d) Translations utilizing the full loss objective, demonstrating that the residual and swap losses are crucial for stabilizing the latent space and preserving semantic structure.
@@ 5.6.1 Pairwise Interpolation
We demonstrate direct $K$ head interpolation, by performing the operation $K_{\text{interp}}=(1-\alpha)K_{i}+\alpha K_{j}$ or $v=(1-\alpha)v_{i}+\alpha v_{j}$, for $\alpha\in[0,1].$ Figure 15 shows samples from interpolating between two classes from the AFHQ dataset, trained for multi-class generation in Section 5.3. We can see a smooth semantic transition between all three class pairs. Each interpolation produces coherent intermediate states, validating that independent $K$ heads enable linear composition directly over the K heads or in velocity space.
@@
Figure 15: Pairwise interpolation samples ($512\times 512$ px) from GAF trained with one $J$ and three independent $K$ heads on the AFHQ dataset.
5.6.2 Continuous Cyclic Transport.
Cyclic Endpoint Transport. We quantitatively analyze GAF’s endpoint transport by running three RK4 [4, 15] integration steps on AFHQ dataset (Cat $\to$ Dog $\to$ Wild $\to$ Cat, Dog $\to$ Wild $\to$ Cat $\to$ Dog, and Wild $\to$ Cat $\to$ Dog $\to$ Wild) and measuring LPIPS [48], pixel MSE (post-VAE), and latent MSE (pre-VAE) between the initial and final states.
We use the cross-modality transport operator via $J$ defined in Section 3.3. Each cycle (e.g., Cat $\to$ Dog, Dog $\to$ Wild, Wild $\to$ Cat) is implemented as the encode–decode path: $v_{i}=K_{i}-J$, using the following method:
(A) Decode from the shared noise anchor $z_{y}$ to class $i$ with $v_{i}$ (forward in time),
(B) Encode back to $J$ with $v_{i}$ (reverse in time), and then,
(C) Decode to the next class $j$ with $v_{i\to j}$.
Figure 16: Cyclic Transport Analysis. We visualize the cyclic transport maps across classes (Top) and their corresponding convergence error rates (Bottom). The trajectories correspond to cyclic transport along the cycle $v_{i\to j\to k\to i}$ across 4 randomly chosen semantic cycles. The error difference between the first and the last image decreases monotonically, reaching machine precision ($\approx 10^{-16}$) at $N=1,000$.
Thus, a full Cat $\to$ Dog $\to$ Wild $\to$ Cat cycle is a sequence of decode–encode legs that always returns through $J$, as in the general cross-modality path $v_{i}\xrightarrow{-(v_{i})}J\xrightarrow{}v_{j}$. The metrics in Figure 16 report the difference between the first and last endpoint of each cycle, both in latent space (pre-VAE) and image space (post-VAE).
Cyclic Velocity Space Interpolation. We demonstrate that GAF enables continuous cyclic transport using cyclic velocity space interpolation, as shown in Figure 17. We use the same multi-class model as in Section 5.3. For a given pair of classes ($i,j$), we define a parametric family of interpolated data spaces or velocity fields:
$K_{i\to j}=(1-\alpha)K_{i}+\alpha K_{j},\quad\alpha\in[0,1],$
$v_{i\to j}=(1-\alpha)v_{i}+\alpha v_{j},\quad\alpha\in[0,1],$
where $v_{i}=K_{i}-J$ and $v_{j}=K_{j}-J$ are the class-specific velocity fields. The full cyclic interpolation between the three classes $i,j$, and $k$ is computed as:
$K_{i}\xrightarrow{J}K_{j}\xrightarrow{J}K_{k}\xrightarrow{J}K_{i}$
$v_{i}\xrightarrow{-v_{i}}J\xrightarrow{}v_{i\to j}\xrightarrow{-v_{i\to j}}J\xrightarrow{}v_{j\to k}\xrightarrow{-v_{j\to k}}J\xrightarrow{}v_{i}$
We sample $\alpha$ at $10$ uniformly spaced values in $[0,1]$ and, for each $\alpha$, refine the endpoint or integrate the velocity field by recovering the noise endpoint from the previous data endpoint using IER or the same ODE solver. Class $i$ is the source class and $j$ is the target class; the interpolated velocity field $v_{i\to j}$ moves from $i$-manifold toward the $j$-manifold while changing only the velocity field, not the underlying integration scheme or random seed. The velocity field $v_{i\to j}$ maintains semantic coherence throughout the transition across classes.
Across all three cycles and $3,000$ random $z_{0}$ samples, we observe an average latent LPIPS of $\approx$ $1.3\times 10^{-16}$ between the initial and the final images in each cycle ($N=1,000$). This indicates GAF’s transport algebra is deterministically reversible. The near-perfect closure validates that independent $K$ heads form a consistent algebraic structure where cyclic compositions return exactly to their starting point.
@@
Figure 17: Continuous cyclic interpolation transport preserving identity across three classes. Each row shows a complete cycle starting from one class (left) to two interpolated classes, and finally returning to the initial class. Distinctive features (coloring, facial marking) are semantically preserved throughout each cycle while class-specific structure transforms.
5.6.3 Spatial Velocity Interpolation
Figure 1819, and 20 shows that spatial composition can be performed either in endpoint space using $K$ heads or in velocity space using $v$. Using region masks ($m$), we combine class-specific endpoint heads directly (e.g., ($K_{\text{cat}},K_{\text{dog}},K_{\text{wild}}$)) to assign different semantics to different facial regions, while keeping the remaining region fixed. For example, we can assign masks to target specific region by dividing the area into $N\in\mathbb{R}$ segments horizontally, diagonally, radially, or in any random configuration, as shown in Figure 20. We can also assign masks to regions such as ears (E), eyes (I), mouth (M), nose (N), and rest (R) (the remaining region not covered by the masks), as shown in Figure 18, to composite features that are directed by the IER or velocity space integration. This produces coherent localized edits that preserve global identity in non-targeted regions, indicating that endpoint composition is a native control mechanism in GAF. Endpoint mixing offers a direct coordinate-level generation (see Figure 19), while velocity mixing uses the ODE generation (see Figure 18 and 20).
Figure 18: Spatial composition on AFHQ using masked velocity blending.
Figure 19: Spatial composition on AFHQ using masked $K$ head blending (8-step IER).
Figure 20: Spatial composition on ImageNet using masked velocity blending.
@@ 5.6.4 Barycentric Transport
We extend pairwise interpolation to a three-way composition using barycentric coordinates. The blended velocity field is $v_{\text{blend}}=\alpha v_{i}+\beta v_{j}+\gamma v_{k}$, where $\alpha+\beta+\gamma=1$. We use square-to-simplex mapping with $\alpha=u(1-u),~\beta=(1-u)(1-v),$ and $\gamma=v$ to cover the full compositional space.
@@
Figure 21 shows three examples of three-way compositions, namely dog-cat-wild, cat-dog-wild, and wild-dog-cat. It shows all weighted combinations, with corners representing pure classes and the interior points showing multi-class blends.
5.7 Comparison with Rectified Flow
GAF learns endpoint operators $J$ and $K$ and samples via IER, while Rectified Flow learns a velocity field $v_{\theta}(\mathbf{x_{t}},t)$ and requires ODE integration for sampling. GAF has an equivalent velocity field v=KJv=K-J. Under ODE sampling, both models exhibit comparable compositional capacity.
Figure 22: GAF comparison with Rectified Flow under identical compositions. The first three images in the grid are the base images, while the fourth is the weighted blend.
Figures 22a, 22b, and 22c illustrate endpoint composition with IER and its qualitative comparison to ODE-based sampling. IER operates directly on the predicted coordinates without solving an ODE, a capability Rectified Flow’s velocity-only formalism cannot support. Because GAF exposes its two endpoints, $J$ and $K$, it also enables Transport Algebra operations directly in endpoint space and native bidirectional sampling, using $K$ for forward generation and $J$ for reverse inversion. Neither capability is present in velocity-only formalism. For example, one can compose endpoints as $0.5K_{\text{dog}}+0.3K_{\text{cat}}+0.2K_{\text{wild}}$ to generate a sample with mixed dog, cat, and wild features. Rectified Flow approximates this by constructing composite velocity field but must route through ODE integration. In contrast, GAF supports endpoint composition via IER, velocity composition via ODEs, and mid-trajectory switching between the two.
@@ 6 Conclusion
We introduced Generative Anchored Fields (GAF), a model that reframes generative modeling from direct trajectory prediction to endpoint coordinate learning. By learning endpoints $J$ and $K$ and deriving an emergent velocity field as v=KJv=K-J, GAF supports three inference methods within a single framework: endpoint iterative refinement (IER), ODE integration, and hybrid switching between them. Since GAF exposes its two endpoints, generation can be treated as algebra over learned transport coordinates. We demonstrated that GAF achieves FID $7.51$ on ImageNet $256\times 256$ without classifier-free guidance. GAF’s factorization into a trunk, $J$, and $K$ yeilds inherent modularity such that each component can be trained, frozen, or swapped independently. Since GAF separates representation learning (trunk) and transport parametrization (JJ/$K$ heads), pretrained vision backbones can serve as GAF’s shared trunk, inheriting Transport Algebra through $J/K$ head attachment without the need to train from scratch.